OpenStudy (anonymous):

Solve the equation 2x^2+24=0, Write the solution in imaginary form

OpenStudy (anonymous):

\[\pm \sqrt{-12} = ...\]

OpenStudy (anonymous):

Does not show that answer under the question ? :o

OpenStudy (anonymous):

That isn't the answer.

OpenStudy (anonymous):

You need to simplify it...

OpenStudy (anonymous):

and how will i do that ? :o,

OpenStudy (anonymous):

\[\sqrt{-12} = \sqrt{-1}\sqrt{12} = \ ...\] You know what root(-1) is (i hope!) if you are doing imaginary numbers, and sqrt 12 can be simplified into the form\[a\sqrt{b}\]

OpenStudy (anonymous):

Yes i do, thanks for the helpp ^-^